# MAP: Unleashing Hybrid Mamba-Transformer Vision Backbone’s Potential with Masked Autoregressive Pretraining

### v1

@@ Abstract
Mamba has achieved significant advantages in long-context modeling and autoregressive tasks, but its scalability with large parameters remains a major limitation in vision applications. pretraining is a widely used strategy to enhance backbone model performance. Although the success of Masked Autoencoder in Transformer pretraining is well recognized, it does not significantly improve Mamba’s visual learning performance. We found that using the correct autoregressive pretraining can significantly boost the performance of the Mamba architecture. Based on this analysis, we propose Masked Autoregressive Pretraining (MAP) to pretrain a hybrid Mamba-Transformer vision backbone network. This strategy combines the strengths of both MAE and Autoregressive pretraining, improving the performance of Mamba and Transformer modules within a unified paradigm. Additionally, in terms of integrating Mamba and Transformer modules, we empirically found that inserting Transformer layers at regular intervals within Mamba layers can significantly enhance downstream task performance. Experimental results show that both the pure Mamba architecture and the hybrid Mamba-Transformer vision backbone network pretrained with MAP significantly outperform other pretraining strategies, achieving state-of-the-art performance. We validate the effectiveness of the method on both 2D and 3D datasets and provide detailed ablation studies to support the design choices for each component.
Figure 1: We propose Masked Autoregressive Pretraining(MAP) to pretrain the hybrid Mamba-Transformer vision backbones. This strategy combines the strengths of both MAE and Autoregressive, improving the performance of Transformer and Mamba modules within a unified paradigm.
@@ 1 Introduction
The State Space Model(Hamilton, 1994) has demonstrated strong capabilities in long-context language modeling. The recent emergence of the variant framework Mamba(Gu & Dao, 2023) has sparked interest in comparing its abilities with those of Transformers. Due to its linear complexity and selective scanning mechanism, Mamba shows significant advantages in computational efficiency when handling long contexts. However, Mamba-based architectures(Zhu et al., 2024b) are difficult to scale concerning the number of parameters, which poses a major limitation for vision applications. To enhance Mamba-based backbones for vision tasks, there’s a trend of combining Mamba with Transformers to create hybrid backbones(Lieber et al., 2024; Hatamizadeh & Kautz, 2024), leveraging the strengths of both. However, to truly scale up these hybrid vision backbones, a good pretraining strategy is essential for maximizing the combined capabilities of Mamba and Transformer. Our work aims to take the first step in this direction.
Developing an effective pretraining strategy for Mamba-Transformer vision backbones is challenging. Even for purely Mamba-based backbones, pretraining methods are still underexplored, and the optimal approach remains unclear. Additionally, the hybrid structure requires a pretraining strategy compatible with both computation blocks. This is particularly challenging because the State Space Model captures visual features very differently from Transformers.
To address these challenges, we conducted extensive pilot studies and identified three key observations. Firstly, existing popular pretraining strategies for Transformers, such as MAE(He et al., 2022) and Contrastive Learning(CL)(He et al., 2020), do not yield satisfactory results for Mamba-based backbones, highlighting the need for a more suitable method. Secondly, Autoregressive Pretraining(AR)(Ren et al., 2024) can be effective for Mamba-based vision backbones, provided that an appropriate scanning pattern and token masking ratio are employed. Thirdly, pretraining strategies suitable for either Mamba or Transformers may not effectively benefit the other, and hybrid backbones require a tailored approach to address the learning needs of different computation blocks.
@@
Our experiments demonstrate that hybrid Mamba-Transformer models pretrained with MAP outperform other pretraining strategies by a significant margin. MAP with the hybrid Mamba-Transformer and pure Mamba backbone can both achieve impressive results on the ImageNet-1k(Deng et al., 2009a) classification task and other 3D vision tasks(Yi et al., 2016; Wu et al., 2015; Uy et al., 2019b). Furthermore, we tried different hybrid integration strategies for combining Mamba and Transformer layers showing that placing Transformer layers at regular intervals within Mamba layers led to a substantial boost in downstream task performance.
Our contributions are threefold: Firstly, we propose a novel method for pretraining the Hybrid Mamba-Transformer Vision Backbone for the first time, enhancing the performance of hybrid backbones as well as pure Mamba and pure Transformer backbones within a unified paradigm. Secondly, we conduct an in-depth analysis of the key components of Mamba with autoregressive pretraining, revealing that the effectiveness hinges on maintaining consistency between the pretraining order and the Mamba scanning order, along with an appropriate token masking ratio. Thirdly, we demonstrate that our proposed method, MAP, significantly improves the performance of both Mamba-Transformer and pure Mamba backbones across various 2D and 3D datasets.
@@ 2 Related Work
Vision Mambas and Vision Transformers. Vision Mamba(Vim)(Zhu et al., 2024a) is an efficient model for visual representation learning, leveraging bidirectional state space blocks to outperform traditional vision transformers like DeiT in both performance and computational efficiency. The VMamba(Liu et al., 2024) architecture, built using Visual State-Space blocks and 2D Selective Scanning, excels in visual perception tasks by balancing efficiency and accuracy. Autoregressive pretraining(ARM)(Ren et al., 2024) further boosts Vision Mamba’s performance, enabling it to achieve superior accuracy and faster training compared to conventional supervised models. Nevertheless, why autoregression is effective for Vision Mamba and what the key factors are remains an unresolved question. In this paper, we explore the critical design elements behind the success of Mamba’s autoregressive pretraining for the first time. Vision Transformers(ViT)(Dosovitskiy, 2020) adapt transformer architectures to image classification by treating image patches as sequential tokens. Swin Transformer(Liu et al., 2021) introduces a hierarchical design with shifted windows, effectively capturing both local and global information for image recognition. MAE (He et al., 2022) enhances vision transformers through self-supervised learning, where the model reconstructs masked image patches using an encoder-decoder structure, enabling efficient and powerful pretraining for vision tasks. However, the MAE pretraining strategy is not effective for Mamba, which hinders our ability to pretrain the hybrid Mamba-Transformer backbones.
@@
Self-Supervised Visual Representation Learning. Self-Supervised Visual Representation Learning is a machine learning approach that enables the extraction of meaningful visual features from large amounts of unlabeled data. This methodology relies on pretext tasks, which serve as a means to learn representations without the need for explicit labels. GPT-style AR(Han et al., 2021) models predict the next part of an image or sequence given the previous parts, encouraging the model to understand the spatial or temporal dependencies within the data. MAE(He et al., 2022) methods mask out random patches of an input image and train the model to reconstruct these masked regions. This technique encourages the model to learn contextual information and global representations. Contrastive Learning(CL)(He et al., 2020) techniques involve contrasting positive and negative samples to learn discriminative features. It typically involves creating pairs of positive and negative examples and training the model to distinguish between them. However, we found that existing pretraining strategies fail to fully unlock the potential of the hybrid framework, which motivated us to explore a new pretraining paradigm for hybrid Mamba-Transformer backbones.
3 Pilot Study: How to pre-train the visual Mamba backbones?
@@
In this Section, we first conduct experiments to investigate the differences in pretraining strategies for ViT and Vim. The success of the MAE strategy on the ViT architecture is well acknowledged, while the Vim pretraining strategy remains in its early stages. We are interested in determining whether the MAE strategy is equally applicable to Vim or if the AR strategy is more suitable. To explore this, we conduct experiments on the classification task using the ImageNet-1K dataset. The results are shown in Table 1.
@@
Table 1: Pilot Study. We use ViT-B and Vim-B as the default configurations. The AR strategy processes the image tokens in a row-first order, while the MAE operates according to the default settings. For contrastive learning, we only used crop and scale data augmentation and used the MoCov2 for pretraining. All experiments are conducted at a resolution of 224x224. The number of mask tokens for AR is set to 40 tokens (20%). Experiments show that MAE is more suitable for Transformer pretraining, while AR is better suited for Mamba pretraining.
We observe that the MAE strategy significantly enhances the performance of ViT. However, for Vim, the MAE strategy does not yield the expected improvements, while the AR strategy substantially boosts its performance. This indicates that for the ViT architecture, applying the MAE strategy is essential to establish bidirectional associations between tokens, thereby improving performance. In contrast, for Vim, it is more important to model the continuity between preceding and succeeding tokens. Based on this observation, we conducted an in-depth analysis of the various components involved in AR pretraining for Mamba and discovered that consistent autoregression pretraining with scanning order and proper masking ratio is the key to pretraining Mamba.
Relationship between AR and Scanning Order. Since the goal of AR pretraining is to learn a high-quality conditional probability distribution, enabling the model to generate new sequences based on previously generated content, we first explore how the prediction order in auto-regressive models affects the pretraining of Vim. Different prediction orders can significantly impact how the model captures image features and the effectiveness of sequence generation. By adjusting the prediction order, we can gain deeper insights into Vim’s behavior in sequence generation tasks and how to effectively model dependencies between elements in an image. Further analysis of the role of prediction order will help optimize AR pretraining for Vim, exploring how the model can better capture the continuity and relationships of image information under different contextual conditions. We conduct ablation studies on Vim by allowing it to perform both row-first and column-first scanning. We then pretrain it with row-first and column-first AR orders, respectively, to compare their performance. Figure 3 shows different orders for AR pretraining and Mamba scanning.
@@
Table 2: The impact of AR pretraining order on downstream tasks. Vim(R) refers to Vim with row-first scanning. Vim(C) refers to Vim with column-first scanning. AR(R) refers to row-first autoregressive pretraining. AR(C) refers to column-first autoregressive pretraining. The results indicate that the best performance is achieved when the auto-regressive pretraining design aligns with Mamba’s scanning order.
@@
The results are shown in Table 2. We observe that employing a pretraining strategy consistent with the scanning order significantly enhances Vim’s performance. This suggests that when designing pretraining strategies, they should be aligned with the downstream scanning order.
Masking Ratio of Autoregression Pretraining. Since the success of MAE is primarily attributed to the use of an appropriate masking ratio, we are inspired to conduct experiments to verify whether different auto-regressive masking ratios will affect the quality of pretraining. We found that during AR pretraining, masking a certain number of tokens at the end of the sequence is crucial. Masking a single token follows the traditional AR paradigm, while masking $n$ tokens transforms the task into an inpainting problem, as the input and output sequence lengths remain equal. In this context, varying the auto-regressive masking ratios effectively adjusts the inpainting ratio, influencing the model’s predictions beyond just the sequence length.
Our pretraining sequence length was set to 196 tokens, and we masked 1 token (0.5%), 20 tokens (10%), 40 tokens (20%), 60 tokens (30%), 100 tokens (50%), and 140 tokens (70%), respectively, while also recording the results of fine-tuning on downstream tasks. Figure 4 shows the pipeline of AR Pretraining under different mask ratios.
@@
Figure 4: Masking Ratio of Autoregression Pretraining. We showcased the autoregressive training process at various masking ratios. Notably, in autoregressive pretraining, different masking ratios effectively control not only the prediction step size but also the length of the input sequence.
@@
The results shown in Table 3 indicate that a proper masking ratio contributes to training stability, helping to avoid excessive noise interference. In auto-regressive pretraining, as the Masking Ratio increases, the performance of the Mamba improves. This is because a higher Masking Ratio encourages the model to learn more complex and rich feature representations, thereby enhancing its generative ability and adaptability. However, an excessively high Masking Ratio may lead to instability during the training process and result in incomplete information perception. We found there exists a sweet spot around 20% on the ImageNet-1K classification task. In such cases, the model may struggle to make accurate predictions due to a lack of sufficient contextual information, negatively impacting its pretraining effectiveness. Therefore, when designing auto-regressive pretraining tasks, finding an appropriate masking ratio is crucial to strike a balance between performance improvement and training stability.
Given that MAE is more suitable for Transformers while AR is better suited for Mamba, how should we approach the pretraining of a hybrid Mamba-Transformer model? We need a new pretraining strategy that is effective for both Transformers and Mamba to support the pretraining of hybrid models. In the next Section, we will provide a detailed explanation of how to pretrain the hybrid Mamba-Transformer backbones.
4 Masked Autoregressive Pretraining for Hybrid Backbones
Our approach represents a general paradigm applicable to data across various domains, with 2D image data as an example. Our method can be easily extended to large language models (LLMs) and the fields of image video and point cloud video. Our method optimizes the synergy between Mamba and Transformer within a unified framework, allowing both models to fully leverage their strengths. In the Mamba-Transformer hybrid architecture, this approach effectively enhances the cooperation between the models, resulting in significant performance improvements. Specifically, our approach includes a masking strategy, a hybrid Mamba-Transformer encoder, and a Transformer decoder. The hybrid Mamba-Transformer encoder is responsible for mapping the signals into latent space, while the Transformer decoder autoregressively reconstructs the features back into the original image. The following section will introduce the specific design components of the framework. The subsequent experiments in this section are conducted using the base-sized model on the ImageNet-1K dataset.
Masking. Consistent with MAE, we first tokenize the image and then apply random masking to a portion of the tokens. We experimented with different masking strategies, including random, sequential, and diagonal masking. Our experiments show that random masking delivers the best results. We attribute this to the fact that sequential and diagonal masking can hinder the Transformer’s ability to establish contextual relationships. Random masking not only promotes bidirectional modeling for Transformers but also enhances Mamba’s generalization and representation capabilities in sequence modeling. Additionally, we explored the effects of different masking ratios and found that a 50% masking ratio yielded the best results. This conclusion aligns with intuition: while MAE performs optimally on Transformers with a 75% masking ratio, previous experiments showed that AR achieves the best results on Mamba with a 20% ratio. Therefore, a 50% ratio serves as a balanced number, leveraging the strengths of both paradigms.
Masking Design From Scratch Random Masking Sequential Masking Diagonal Masking
Accuracy 83.1 84.9 84.0 83.8
Masking Ratio 0% 25% 50% 75%
Accuracy 83.3 84.5 84.9 84.2
Table 4: Random masking with a 50% masking ratio performs the best.

### v2

@@ Abstract
Hybrid Mamba-Transformer networks have recently garnered broad attention. These networks can leverage the scalability of Transformers while capitalizing on Mamba’s strengths in long-context modeling and computational efficiency. However, the challenge of effectively pretraining such hybrid networks remains an open question. Existing methods, such as Masked Autoencoders (MAE) or autoregressive (AR) pretraining, primarily focus on single-type network architectures. In contrast, pretraining strategies for hybrid architectures must be effective for both Mamba and Transformer components. Based on this, we propose Masked Autoregressive Pretraining (MAP) to pretrain a hybrid Mamba-Transformer vision backbone network. This strategy combines the strengths of both MAE and Autoregressive pretraining, improving the performance of Mamba and Transformer modules within a unified paradigm. Experimental results show that the hybrid Mamba-Transformer vision backbone network pretrained with MAP significantly outperforms other pretraining strategies, achieving state-of-the-art performance. We validate the method’s effectiveness on both 2D and 3D datasets and provide detailed ablation studies to support the design choices for each component. Code and checkpoints are available at https://github.com/yunzeliu/MAP.
@@ 1 Introduction
The Mamba-Transformer backbone[16, 24, 49, 5] has recently attracted widespread attention. It leverages the scalability advantages of Transformers and utilizes Mamba’s[11, 62] strong capabilities in long-context language modeling. However, to scale up Mamba-Transformer vision backbones, an effective pretraining strategy is essential for maximizing the combined capabilities of Mamba and Transformer. Our work aims to take the first step in this direction.
Figure 1: We propose Masked Autoregressive Pretraining to pretrain the hybrid Mamba-Transformer backbone. It demonstrates significant performance improvements on both 2D and 3D tasks.
@@
Developing an effective pretraining strategy for Mamba-Transformer vision backbones is challenging. Although the capability of MAE in pretraining Transformers has been well validated, pretraining methods for Mamba are still underexplored, and the optimal approach remains unclear. Additionally, the hybrid structure requires a pretraining strategy compatible with both computation blocks. This is particularly challenging because the State Space Model captures visual features very differently from Transformers.
To address these challenges, we conducted pilot studies and identified three key observations. Firstly, existing popular pretraining strategies for Transformers, such as MAE[19] and Contrastive Learning[18], do not yield satisfactory results for Mambas. Secondly, Autoregressive Pretraining[41] can be effective for Mamba-based vision backbones with an appropriate scanning pattern and token masking ratio. Thirdly, pretraining strategies suitable for either Mamba or Transformers do not effectively benefit the other.
@@
Notably, since our focus in this paper is on pretraining strategies for hybrid frameworks rather than the hybrid frameworks themselves, we empirically choose a densely mixed hybrid scheme as the default hybrid framework and refer to it as HybridNet. Our experiments demonstrate that HybridNet pretrained with MAP outperforms other pretraining strategies by a significant margin on the 2D and 3D vision tasks. Extensive ablation studies confirm the rationale and effectiveness of each design choice.
Our contributions are threefold as follows:
Firstly, we propose a novel method for pretraining the Hybrid Mamba-Transformer Vision Backbone for the first time, enhancing the performance of hybrid backbones as well as pure Mamba and pure Transformer backbones within a unified paradigm. Secondly, For diagnostic purposes, we also conduct an in-depth analysis of the key components of pure-Mamba with autoregressive pretraining, revealing that the effectiveness hinges on maintaining consistency between the pretraining order and the Mamba scanning order, along with an appropriate token masking ratio. Thirdly, we demonstrate that our proposed method, MAP, significantly improves the performance of both Mamba-Transformer and pure Mamba backbones across various 2D and 3D datasets. Extensive ablation studies confirm the rationale and effectiveness of each design choice.
@@ 2 Related Work
Vision Mambas and Vision Transformers. Vision Mamba(Vim)[63] is an efficient model for visual representation learning, leveraging bidirectional state space blocks to outperform traditional vision transformers like DeiT in both performance and computational efficiency. The VMamba[29] architecture, built using Visual State-Space blocks and 2D Selective Scanning, excels in visual perception tasks by balancing efficiency and accuracy. Autoregressive pretraining(ARM)[41] further boosts Vision Mamba’s performance, enabling it to achieve superior accuracy and faster training compared to conventional supervised models. Nevertheless, why autoregression is effective for Vision Mamba and what the key factors are remains an unresolved question. In this paper, we explore the critical design elements behind the success of Mamba’s autoregressive pretraining for the first time. The key difference between our proposed MAP and ARM[41] lies in our use of random masking in the AR process, where the network decodes the next complete local information based on the masked local information. The masking mechanism enhances the modeling capability of local information and the relevance of context. Vision Transformers(ViT)[8] adapt transformer architectures to image classification by treating image patches as sequential tokens. Swin Transformer[30] introduces a hierarchical design with shifted windows, effectively capturing both local and global information for image recognition. MAE [19] enhances vision transformers through self-supervised learning, where the model reconstructs masked image patches using an encoder-decoder structure, enabling efficient and powerful pretraining for vision tasks. However, the MAE pretraining strategy is not effective for Mamba, which hinders our ability to pretrain the hybrid Mamba-Transformer backbones. Although the hybrid architecture using MAP still struggles to surpass the Transformer using MAE under the same settings, it provides a balance between computational cost and performance, as well as the potential to introduce longer contexts while maintaining scalability. This property is particularly important in the video domain and large language models, and we leave it for future work.
@@
Self-Supervised Visual Representation Learning. Self-Supervised Visual Representation Learning is a machine learning approach that enables the extraction of meaningful visual features from large amounts of unlabeled data. This methodology relies on pretext tasks, which serve as a means to learn representations without the need for explicit labels. GPT-style AR[14] models predict the next part of an image or sequence given the previous parts, encouraging the model to understand the spatial or temporal dependencies within the data. MAE[19] methods mask out random patches of an input image and train the model to reconstruct these masked regions. This technique encourages the model to learn contextual information and global representations. Contrastive Learning(CL)[18] techniques involve contrasting positive and negative samples to learn discriminative features. It typically involves creating pairs of positive and negative examples and training the model to distinguish between them. However, we found that existing pretraining strategies fail to fully unlock the potential of the hybrid framework, which motivated us to explore a new pretraining paradigm for hybrid Mamba-Transformer backbones. In this paper, we propose MAP to pretrain hybrid architectures. On the pure Mamba architecture, our method outperforms AR, due to MAP’s ability to effectively model local features and the associations between local regions. On the pure Transformer architecture, although MAP does not surpass the performance of MAE, it still achieves comparable results. This is because, while local MAE reduces the receptive field, the autoregressive modeling between local regions enhances the ability to capture local relationships, ensuring that MAP maintains strong performance on Transformers.
3 Polit Study
@@
Before exploring the strategies for pretraining hybrid frameworks, we first analyze how to properly pretrain Transformers and Mamba respectively. Analyzing the pretraining strategies for them helps us gain a deeper understanding of the key aspects of training hybrid frameworks. In this Section, we first conduct experiments to investigate the differences in pretraining strategies for ViT and Vim, which are representative of visual Transformers and visual Mamba. The success of the MAE strategy on the ViT architecture is well acknowledged, while the Vim pretraining strategy remains in its early stages. We are interested in determining whether the MAE strategy is equally applicable to Vim or if the AR strategy is more suitable. To explore this, we conduct experiments on the classification task using the ImageNet-1K dataset. The results are shown in Table 1.
@@
Table 1: Pilot Study. We use ViT-B and Vim-B as the default configurations. The AR strategy processes the image tokens in a row-first order, while the MAE operates according to the default settings. For contrastive learning, we only used crop and scale data augmentation and used the MoCov2 for pretraining. All experiments are conducted at a resolution of 224x224. The number of mask tokens for AR is set to 40 tokens (20%).
Experiments show that MAE is more suitable for Transformer pretraining, while AR is better suited for Mamba pretraining. We observe that the MAE strategy significantly enhances the performance of ViT. However, for Vim, the MAE strategy does not yield the expected improvements, while the AR strategy substantially boosts its performance. This indicates that for the ViT, applying the MAE strategy is essential to establish bidirectional associations between tokens, thereby improving performance. In contrast, for Vim, it is more important to model the continuity between preceding and succeeding tokens. Therefore, for the hybrid framework, we intuitively need a strategy that can leverage the advantages of MAE in bidirectional modeling while retaining the strengths of AR in context modeling. Since MAE’s pretraining on Transformers is already well-established, benefiting from its asymmetric network design and high masking ratio, the key to AR’s success on Mamba remains unclear. Therefore, in the following sections, we analyze the critical factors behind AR’s success on Mamba, which will help us design strategies for the hybrid framework later.
We conducted an in-depth analysis and discovered that consistent autoregression pretraining with scanning order and proper masking ratio is the key to pretraining Mamba. This is crucial for diagnosing how to pretrain Mamba and for designing pretraining strategies for the hybrid backbone.
Relationship between AR and Scanning Order. Different prediction orders can significantly impact how the model captures image features and the effectiveness of sequence generation. The analysis of the role of prediction order will help optimize AR pretraining for Vim, exploring how the model can better capture the continuity and relationships of image information under different contextual conditions. We conduct ablation studies on Vim by allowing it to perform both row-first and column-first scanning. We then pretrain it with row-first and column-first AR orders, respectively, to compare their performance. Figure 3 shows different orders for AR pretraining and Mamba scanning.
@@
Table 2: The impact of AR order on downstream tasks. Vim(R) refers to Vim with row-first scanning. Vim(C) refers to Vim with column-first scanning. AR(R) refers to row-first autoregressive pretraining. AR(C) refers to column-first autoregressive pretraining. The results indicate that the best performance is achieved when the AR pretraining design aligns with Mamba’s scanning order.
@@
The results are shown in Table 2. We observe that employing a pretraining strategy consistent with the scanning order significantly enhances Vim’s performance. This suggests that when designing pretraining strategies, they should be aligned with the downstream scanning order. This conclusion is also implicitly confirmed in ARM[41], as it uses cluster-based AR to pretrain cross-scanned Mamba, which is essentially a combination of row-first and column-first approaches. Unlike ARM, we experimentally verified this and provided comprehensive experimental results.
Masking Ratio of Autoregression Pretraining. Masking a single token follows the traditional AR paradigm, while masking $n$ tokens transforms the task into an inpainting problem, as the input and output sequence lengths remain equal. In this context, varying the auto-regressive masking ratios effectively adjusts the inpainting ratio, influencing the model’s predictions beyond just the sequence length.
@@
The results are shown in Table 3. In auto-regressive pretraining, as the Masking Ratio increases, the performance of the Mamba improves. This is because a higher Masking Ratio encourages the model to learn more complex and rich feature representations, thereby enhancing its generative ability and adaptability. However, an excessively high Masking Ratio may lead to instability during the training process. In such cases, the model may struggle to make accurate predictions due to a lack of sufficient contextual information, negatively impacting its pretraining effectiveness. Therefore, when designing auto-regressive pretraining tasks, finding an appropriate masking ratio is crucial to strike a balance between performance improvement and training stability.
Now we have the following three conclusions to serve as a reference for designing the pretraining of hybrid backbone:
  • MAE is more suitable for Transformers, while AR is better suited for Mamba.
  • For MAE pretraining of Transformers, an asymmetric structure, and an appropriate masking ratio are important.
  • For AR pretraining of Mamba, an appropriate AR order and masking ratio are important.
Based on the above conclusions, we propose Masked Autoregressive Pretraining (MAP) to pretrain hybrid frameworks. MAP leverages AR to model the relationships between local regions and local MAE to model the internal features of local regions. It also employs a masking strategy that is crucial for both MAE and AR to enhance the network’s representation capabilities. This strategy combines the strengths of MAE in local feature modeling, which is suitable for pretraining Transformers, and the strengths of AR in context modeling, which is suitable for pretraining Mamba. Effectively inheriting the advantages of both strategies, MAP improves the pretraining performance of hybrid architectures within a unified framework. In the following section, we will provide a detailed introduction to the MAP pretraining strategy.
4 Method